# Point Cloud Streaming with Latency-Driven Implicit Adaptation using MoQ

### v1

@@ IV-B Results
TABLE II: Average change in throughput (%) compared to the baseline 50 ms timeout.
We provide the average change in throughput across our various test conditions in Tab. II. A key result is that throughput generally correlates with the latency allowance. This is evident at a low bandwidth (300 Mbps) and moderate timeout (500 ms). Importantly, these throughput gains correspond to improvements in the PCQM quality, as shown in Tab. I. In Fig. 2, we illustrate the rate of “stall” scenarios, which we define as frames for which we receive zero representations within the delivery timeout. Although the use of only 5 Tracks yields higher average visual quality under low bandwidth, it also induces severe stall rates at low timeouts. We emphasize that the quality impact of a playback stall is subjective and should be an area of future investigation.
@@
Under the 10-Track configuration, we observe an improvement when shifting from 5 FPG to 30 FPG. Specifically, we see fewer stall events, higher throughput, and better visual quality. For example, with 600 Mbps bandwidth, we observe an average 90.8% increase in throughput and a 0.0024-point improvement in the PCQM metric when increasing the latency allowance from 50 ms to 500 ms. This result is somewhat surprising, as we expected that the faster recovery with smaller Groups would yield higher throughput under network congestion. In practice, the larger Group size functions to better mitigate congestion as a Track waits longer to begin sending again once a Group is dropped. That is, we better avoid overwhelming the QUIC congestion controller by not sending less frequent bursts of data.
Whereas low-latency HAS adaptation simply aims to minimize the end-to-end latency, our system unlocks a unique trade-off between latency and quality. If users can tolerate higher latency, they will receive more sampled representations and experience higher visual quality.
V Conclusion and Future Work
@@
While Random Uniform Sampling offers the low latency necessary for immersive conferencing, we believe that content-adaptive partitioning schemes are a promising direction to derive representations of varying importance. This can be elegantly paired with prioritization mechanisms in MoQ, allowing us to give more precedence to representations deemed more important during transmission. For example, human faces should have a higher priority than their clothing for teleconferencing applications.
Finally, there is significant room to explore the impact of implicit adaptation on the user’s quality of experience (QoE). Our PCQM measurements indicate a predictable latency-distortion trade-off for individual frames, but we do not yet quantify the impact of potentially drastic changes in the received quality or stall scenarios when no representations are received for a given frame. Subjective user studies will be necessary to evaluate and mitigate these factors. From there, we can better determine the optimal encoder and sampling configurations to maximize QoE.

### v2

@@ 4.2. Results
Table 2. Average change in throughput (%) compared to the baseline 50 ms timeout.
4.2.2. Stall rate.
In Fig. 2, we illustrate the rate of “stall” scenarios, which we define as the percentage of frames for which we receive zero representations within the delivery timeout. Using only 5 Tracks results in severe stall rates at short delivery timeouts under low bandwidth, since each Track carries a high data burden. At the same time, when frames are successfully delivered, The 5-Track configurations yield slightly higher per-frame reconstruction quality under low bandwidth (300 Mbps), as reflected by the 1-PCQM scores in Tab. 1, which excludes stall frames. In contrast, configurations with 10 Tracks exhibit significantly lower stall rates across all bandwidths and timeout settings, while maintaining high 1-PCQM scores for the non-stall frames. We note that the Quality of Experience (QoE) impact of a playback stall is subjective and should be an area of future investigation.
@@
Under the 10-Track configuration, we observe an improvement when shifting from 5 FPG to 30 FPG. Specifically, we see fewer stall events, higher throughput, and better visual quality. For example, with 600 Mbps bandwidth, we observe an average 90.8% increase in throughput (Tab. 2) and a 0.0024-point improvement in the PCQM metric (Tab. 1) when increasing the latency allowance from 50 ms to 500 ms. The larger Group size mitigates QUIC congestion through increased spacing between bursts of data.
4.2.3. DASH comparison.
This section compares our MoQ-based system with the four DASH baseline ABR algorithms provided in VV-DASH (Heidarirad and Wang, 2025). These include the default “dash” algorithm, as well as bandwidth-based, buffer-based, and hybrid schemes, which adapt video quality based on network throughput, buffer occupancy, or a combination of both.
Throughput comparison. Fig. 3 compares the average throughput between our MoQ-based system and the DASH baseline across our three bandwidth settings and all encoder/packaging configurations. Under high congestion, the Bandwidth and Buffer ABR algorithms perform best for DASH, exceeded only by our 500 ms and 1500 ms delivery timeouts under MoQ. Whereas DASH has higher throughput for smaller FPG, since it can switch quality level more frequently, MoQ exhibits its best average throughput and variance for the configuration with 10 Tracks and 30 FPG. Under this configuration, MoQ has similar throughput (and expected PCQM) to the best DASH ABR, even when we set an extremely low delivery timeout of 50 ms. Meanwhile, we expect that a similarly-configured LL-DASH (Bentaleb et al., 2022) implementation would exhibit worse throughput (and quality) than our experiments show, since it has a similar trade-off between quality and latency.
Dynamic adaptation. Fig. 4 illustrates the change in throughput over time for a representative example using the 10 Track, 30 FPG configuration. For readability, we show the number of received point cloud representations (via Tracks or by the selected DASH quality level), the high-performing Bandwidth ABR algorithm, and two MoQ delivery timeouts. We observe that when network conditions are unknown at the beginning of the stream session, the explicit client-side ABR switching of DASH has a slow warm-up period, gradually increasing its chosen quality. Meanwhile, our MoQ system delegates implicit adaptation to the transport layer, attempting to send as much data as possible, allowing for higher initial throughput. However, this also brings with it higher variance in throughput, sometimes leading to stall frames as in Fig. 4(a).
5. Content-Aware Streaming PoC
We demonstrate a content-aware streaming proof-of-concept (PoC) that leverages QUIC stream priorities to improve visual quality under our implicit adaptation scheme. Instead of Random Uniform Sampling (Sec. 3.1), we variably sample points based on their membership in detected classes of interest. We trained a PointNet model (Charles et al., 2017) with the EgoBody dataset (Zhang et al., 2022) to perform binary semantic segmentation of human head and non-head classes. This trained model is applied to the MVUB dataset  (Loop et al., 2016) in real time with a GTX 1080 Ti GPU. We then weight our random sampling such that the MoQ Tracks with the highest QUIC priorities contain approximately 50% “head” points and 50% “non-head” points.
Subject to reasonable segmentation accuracy, we find that our MoQ scheme implicitly emphasizes the transmission of the salient “head” points during periods of network congestion. We provide a visual example in Fig. 5, with 2D-rendered frames. We can see that the content-aware sampling yields far fewer holes in the facial region, which is crucial for subjective quality. This yields a 10.1 improvement in average VMAF score under moderate network congestion [^2] We provide an anonymous video example for the reviewers here: https://zenodo.org/records/18203126. This result demonstrates a unique use case to combine MoQ delivery timeouts with deliberate QUIC stream prioritization for content-aware streaming, which may improve QoE for immersive teleconferencing applications.
(a) Random, VMAF: 68.9
(b) Salient, VMAF: 79.0
Figure 5. Impact of content-aware sampling on the andrew9 sequence (Loop et al., 2016), with higher transport priority given to points detected in the head region. Configuration: 10 Tracks, 30 FPG, 600 Mbps bandwidth, 100 ms timeout.
6. Conclusion and Future Work
@@
While Random Uniform Sampling offers the low latency necessary for immersive conferencing, we demonstrated that content-aware partitioning schemes are a promising direction to improve QoE. Further work should explore more advanced models with additional content classes. The application of scalable coding strategies is an alternative future direction, but requires future exploration on the robustness of MoQ/QUIC prioritization mechanisms.
Finally, there is significant room to explore the impact of implicit adaptation on the user’s QoE. Our PCQM measurements indicate a predictable latency-distortion trade-off for individual frames, but we do not yet quantify the impact of potentially drastic changes in the received quality or stall scenarios when no representations are received for a given frame. Subjective user studies will be necessary to evaluate and mitigate these factors. From there, we can better determine the optimal encoder and sampling configurations to maximize QoE.